### v1

# Counting Hallucinations in Diffusion Models

Abstract
Diffusion probabilistic models (DPMs) have demonstrated remarkable progress in generative tasks, such as image and video synthesis. However, they still often produce hallucinated samples (hallucinations) that conflict with real-world knowledge, such as generating an implausible duplicate cup floating beside another cup. Despite their prevalence, the lack of feasible methodologies for systematically quantifying such hallucinations hinders progress in addressing this challenge and obscures potential pathways for designing next-generation generative models under factual constraints. In this work, we bridge this gap by focusing on a specific form of hallucination, which we term counting hallucination, referring to the generation of an incorrect number of instances or structured objects, such as a hand image with six fingers, despite such patterns being absent from the training data. To this end, we construct a dataset suite CountHalluSet, with well-defined counting criteria, comprising ToyShape, SimObject, and RealHand. Using these datasets, we develop a standardized evaluation protocol for quantifying counting hallucinations, and systematically examine how different sampling conditions in DPMs, including solver type, ODE solver order, sampling steps, and initial noise, affect counting hallucination levels. Furthermore, we analyze their correlation with common evaluation metrics such as Fréchet Inception Distance (FID), revealing this widely used image quality metric fail to capture counting hallucinations consistently. This work aims to take the first step toward systematically quantifying hallucinations in diffusion models and offer new insights into the investigation of hallucination phenomena in image generation.
1 Introduction
Diffusion probabilistic models (DPMs) (Ho et al., 2020; Dhariwal & Nichol, 2021; Rombach et al., 2022), commonly referred to as diffusion model, have achieved considerable success in generative modeling, producing high-quality samples across various tasks such as text-to-image generation (Ho & Salimans, 2021; Rombach et al., 2022), audio synthesis (Kong et al., 2021), video generation (Ho et al., 2022), 3D object generation (Luo & Hu, 2021; Vahdat et al., 2022), and protein structure prediction (Jumper et al., 2021; Abramson et al., 2024). Despite these impressive advancements, diffusion models still struggle to faithfully capturing factual properties from training data, often generating outputs that violate consistency with real-world knowledge. A notable example is the generation of objects with an incorrect number of spatially intricate components, such as extra human fingers (Narasimhaswamy et al., 2024) or additional animal limbs (Borji, 2023).
Recent studies have increasingly focused on identifying and analyzing common failure modes in text-guided diffusion models. For instance, Borji (2023) systematically concluded several qualitative shortcomings in text-to-image models. Additionally, Liu et al. (2024) proposed an adversarial search method to uncover undesirable behaviors and failure cases in text-to-image generation, while Wu et al. (2023) employed an iterative approach to discover and mitigate spurious correlations.
More recently, hallucination, a specific failure mode in diffusion models, has garnered attention in the research community (Aithal et al., 2024). Unlike previously discussed text‑prompt‑based failure modes, which can be evaluated by measuring the alignment between the input conditions and outputs (e.g., CLIP score (Hessel et al., 2021) and VQA score (Huang et al., 2025; Zarei et al., 2024)), or more obvious failure modes such as severe distortions, blur, and artifacts, hallucinated samples (hallucinations) often appear visually plausible, and emerge independently of the conditioning signal, manifesting as subtle violations of factual consistency with respect to the training data, such as incorrect object counts, geometries that appear plausible but are physically impossible, or other breaches of physical laws. These fine-grained distinctions between valid samples and hallucinations present a critical challenge for current AI systems, which often fail to detect them reliably. For example, even state-of-the-art large vision-language models like GPT-4o still struggle with basic tasks like accurately counting the number of fingers in a clear hand image. To this end, establishing a standardized and well-defined protocol for reliably quantifying hallucinated samples is essential, as it provides a foundation for outlining pathways toward designing next-generation generative models with strict factual consistency, an ability that is critical for their role as world models (Ha & Schmidhuber, 2018; Ding et al., 2024; Bruce et al., 2024).
Figure 1: Example comparison between the proposed CountHalluSet used for training (left) and the corresponding counting hallucinations generated by the model (right). Each dataset contains different objects with specific counting criteria. For example, in ToyShape and SimObject, each image contains at most one instance per category and at least one instance overall, whereas in RealHand, each image contains exactly five fingers. Counting hallucinations refer to generated images that violate the dataset’s counting rules, such as two apples in an image or an image containing no objects but a normal background, even though these patterns never appear in the SimObject dataset.
Among the various common types of hallucinated elements, such as counts, colors, spatial relationship, geometry, and physical realism, counting errors are particularly suitable for systematic study because they are not only easy to identify but also straightforward to quantify. This motivate us to begin by defining a specific form of hallucination termed counting hallucination. In this setting, the number of objects in training data is explicitly predetermined. Consequently, a generated image is deemed hallucinated whenever its object counts deviate from these predefined values. For example, an image depicting a hand with six fingers is classified as a counting hallucination, since such a pattern never occurs in the RealHand dataset, as shown in the last row of Fig. 1. To facilitate the study of such phenomena, we construct a dataset suite, CountHalluSet, consisting of three datasets that span a spectrum of morphological complexity for the countable objects: ToyShape (triangle, square, pentagon), SimObject (mug, apple, clock), and RealHand (finger). Figure 1 illustrates representative examples from these datasets alongside corresponding counting hallucinations.
In our quantification protocol for counting hallucinations, a critical step is the accurate identification and quantification of objects. To this end, we employ counting models tailored to each dataset. Leveraging these datasets and counting models, we examine how various denoising conditions (e.g., sampling steps, initial noise, the order of ODE solvers, and solver type) within diffusion models affect the counting hallucination rate. Besides, we conduct correlation analysis between counting hallucination rate and Fréchet Inception Distance (FID).
Several significant findings or contributions include:
  • Commonly applied numerical strategies in diffusion models, such as increasing sampling steps from 25 to 100, can help mitigate counting hallucinations in synthetic datasets but exacerbates them in the RealHand dataset.
  • The correlations between counting hallucination rate and FID are dataset-dependent and solver-dependent, rather than intrinsic.
  • Based on one of our findings, we propose a simple but effective method, termed joint-diffusion models, which can significantly reduce counting-based hallucinations and non-counting failures in the RealHand dataset.
2 Related Work
2.1 Hallucinations in Diffusion Models
Inspired by the growing attention to hallucination in large language models (LLMs) (Kadavath et al., 2022; Ji et al., 2023; Huang et al., 2023), the study of hallucinations in unconditional diffusion models has also garnered increasing interest. In particular, Aithal et al. (2024) proposed a qualitative hypothesis suggesting that hallucinated samples may arise from data interpolating between distinct modes. Despite this insight into potential causes, a systematic quantitative analysis of hallucinations in diffusion models remains lacking.
This motivates us to facilitate the accurate identification and quantification of hallucinations in diffusion-generated outputs. Such an investigation can deepen our understanding of undesirable behaviors exhibited by these models, and clarify how the magnitude of denoising errors, such as variations in sampling steps or the order of ODE solvers, affects the severity of hallucinations.
2.2 Quantitative Metrics in Image Generation
The Visual Question Answering (VQA) score (Huang et al., 2025; Zarei et al., 2024) and the CLIP score (Hessel et al., 2021) are two widely used metrics for evaluating the performance of text‑guided generated images. However, such evaluation methods that rely on language‑based models are inherently limited when applied to assess hallucinated images. One line of evidence is the statistical lower bound on the propensity of pretrained language models to hallucinate certain fact types, which constrains the upper bound of detection accuracy based solely on language models (Kalai & Vempala, 2024). Another indication is the persistent performance gap between human annotators and LLMs in classifying hallucinated textual content (Lin et al., 2022; Li et al., 2023).
In addition, several widely used quantitative evaluation metrics, such as FID (Heusel et al., 2017), IS (Salimans et al., 2016), and Precision & Recall (Sajjadi et al., 2018), are commonly employed in image generation tasks. These metrics rely on the feature embeddings extracted from a pretrained Inception-v3 model (Szegedy et al., 2016). However, whether these metrics can capture the severity of hallucinations in generated datasets remains an open question. In this work, we aim to address this gap from the perspective of counting hallucinations.
3 Preliminary: Diffusion Probabilistic Models
3.1 Forward (Diffusion) Process
Forward Process. Given the observed data ${\bm{x}}_{0}\sim q_{0}({\bm{x}}_{0})$, Diffusion Probabilistic Models (DPMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020) introduce a forward process (a.k.a, diffusion process) that gradually perturbs the data by adding Gaussian noise over $T$ discrete timesteps. At at each timestep $t$, the marginal distribution of the noisy sample ${\bm{x}}_{t}$ conditioned on the original data ${\bm{x}}_{0}$ follows:
$q_{0:t}({\bm{x}}_{t}|{\bm{x}}_{0})={\mathcal{N}}({\bm{x}}_{t}\mid\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0},(1-\bar{\alpha}_{t}){\bm{I}}),\quad\alpha_{t}:=1-\beta_{t},\;\bar{\alpha}_{t}:=\prod_{j=1}^{t}\alpha_{j},$ (1)
where $t\in\{0,1,\ldots,T\}$ and $\{\beta_{i}\}_{i=1}^{T}$ is a pre-defined noise schedule dependent on $t$, commonly chosen to be linear (Ho et al., 2020) or cosine (Nichol & Dhariwal, 2021) schedules. The factor $\bar{\alpha}_{t}$ controls the signal-to-noise ratio at step $t$: $\bar{\alpha}_{0}=1$ (no noise) and $\bar{\alpha}_{T}\approx 0$ (nearly pure noise). An equivalent formulation of Eq. (1) is:
${\bm{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+(1-\bar{\alpha}_{t}){\bm{\epsilon}},\quad{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\bm{I}}).$ (2)
As $T$ becomes sufficiently large, the marginal distribution $q_{0:T}({\bm{x}}_{T}|{\bm{x}}_{0})$ approaches a standard Gaussian distribution ${\mathcal{N}}(\bm{0},{\bm{I}})$. In widely-used DPMs (Ho et al., 2020; Nichol & Dhariwal, 2021), $T$ is typically set to 1000.
3.2 Reverse (Denoising) Process.
Training Objective. The reverse process (a.k.a., denoising process) of DPMs aims to iteratively recover the observed data based on the diffused data ${\bm{x}}_{T}$. Specifically, DPMs attempt to predict the noise $\bm{\epsilon}$} from the data ${\bm{x}}_{t}$ by using a neural network ${\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)$, known as noise prediction model. The parameter $\theta$ is optimized by minimizing the following objective (Ho et al., 2020; Song et al., 2022):
$\min_{\theta}\mathbb{E}_{t,{\bm{x}}_{0},{\bm{\epsilon}}}[\pi(t)\|{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)-{\bm{\epsilon}}\|_{2}^{2}],$ (3)
where $\pi(t)>0$ is a weighting function.
Ancestral Sampling. Given noisy data ${\bm{x}}_{T}$, samples from the original distribution $q_{0}({\bm{x}}_{0})$ can be obtained by reversing the diffusion process with the same number of steps, commonly referred to as ancestral sampling:
$p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t})={\mathcal{N}}({\bm{x}}_{t-1}\mid\frac{1}{\sqrt{1-\beta_{t}}}({\bm{x}}_{t}+\beta_{t}\nabla_{{\bm{x}}_{t}}\,log\,q_{t}({\bm{x}}_{t})),\beta_{t}{\bm{I}}),$ (4)
where $\nabla_{{\bm{x}}_{t}}\,log\,q_{t}({\bm{x}}_{t})$ is the only unknown term, referred as score function (Song et al., 2022).
In practice, DPMs approximate the scaled score function $\sqrt{(1-\bar{\alpha}_{t})}\nabla_{{\bm{x}}_{t}}\,log\,q_{t}({\bm{x}}_{t})$ via the noise prediction model. Therefore, an equivalent formulation of Eq. (3) is:
${\bm{x}}_{t-1}={\bm{\mu}}_{\theta}({\bm{x}}_{t},t)+\sqrt{\beta_{t}}{\bm{\epsilon}},\quad{\bm{\mu}}_{\theta}({\bm{x}}_{t},t)=\frac{1}{\sqrt{1-\beta_{t}}}{\bm{x}}_{t}-\frac{\beta_{t}}{\sqrt{1-\beta_{t}}}\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t).$ (5)
ODE-Based Accelerated Sampling. The ancestral sampling in Eq. (4) contains a stochastic term (Kloeden et al., 1992), while there is an associated deterministic process, which can be formulated as an ordinary differential equation (ODE) (a.k.a, probability flow ODE (Song et al., 2022)):
$\frac{\text{d}{\bm{x}}(\tau)}{\text{d}\tau}=f(\tau){\bm{x}}(\tau)+h({\bm{x}}(\tau),\tau),\quad f(\tau)=-\frac{1}{2}\beta(\tau),\;h({\bm{x}}(\tau),\tau)=-\frac{1}{2}\beta(\tau)\nabla_{{\bm{x}}(\tau)}\,log\,q_{\tau}({\bm{x}}(\tau)),$
where $\tau\in[T,0]$. Let ${\bm{x}}(\tau)$ denote the ideal trajectory that satisfies this ODE with the true score function $h({\bm{x}}(\tau),\tau)$ and starts from an ideal diffused data ${\bm{x}}(T)$ drawn from the true diffused distribution $q_{T}({\bm{x}}(T))$. Set the discrete time steps as $\tau_{k}=T-k\Delta t$ for $k=0,1,\ldots,N$, where $\tau_{N}=0$ Applying the variation‐of‐constants formula for the exact solver over one step:
${\bm{x}}(\tau_{k+1})={\mathcal{G}}(\tau_{k+1},\tau_{k}){\bm{x}}(\tau_{k})+\int_{\tau_{k}}^{\tau_{k+1}}{\mathcal{G}}(\tau_{k+1},u)h({\bm{x}}(u),u)\text{d}u$ (7)
where ${\mathcal{G}}(t_{2},t_{1})=e^{\int_{t_{1}}^{t_{2}}f(v)\text{d}v}$ is the state transition operator for the linear part $\frac{\text{d}{\bm{z}}}{\text{d}\tau}=f(\tau){\bm{z}}$. $\Delta t>0$ denotes the step size, allowing ODE-based solvers to accelerate the sampling process with larger step sizes compared to ancestral sampling. The total number of sampling steps is given by $T/\Delta t$. Since the linear part ${\mathcal{G}}(t_{2},t_{1})$ can be computed in closed form, most accelerated ODE-based samplers (Song et al., 2021; Liu et al., 2022; Zhang & Chen, 2023; Lu et al., 2022; 2023) devote their numerical efforts, such as higher solver orders and more sampling steps, to accurately approximating the nonlinear integral term involving $h$. More theoretical analysis can be found in Appendix A.1 to A.3.
4 Formal Definition, Datasets, and Evaluation Protocol for Counting Hallucinations
In this section, we describe the methodology and datasets for quantifying counting hallucinations. We first define counting criteria and the associated counting hallucinations, then introduce the datasets in the CountHalluSet suite. For each dataset, we specify its counting criteria and corresponding counting hallucinations, and outline the standardized protocol used for their quantification.
4.1 Counting Hallucinations: Deviations from the Counting Criteria
In this section, we provide a general definition of counting criteria within training datasets and define counting hallucinations as deviations from these predefined constraints.
Counting criteria in a training dataset. Let ${\mathcal{D}}_{\mathrm{\text{ref}}}$ be a training dataset. For any reference sample ${\bm{x}}\in{\mathcal{D}}_{\mathrm{\text{ref}}}$, the number of objects of category $c$ in $\bm{x}$}, $N_{c}({\bm{x}})$, are given by a predefined set ${\mathcal{S}}_{c}\subseteq\mathbb{Z}_{\geq 0}$ (e.g., $\{0,1\}$ or $\{1,2,3\}$), i.e., $N_{c}({\bm{x}})\in{\mathcal{S}}_{c}$. We further require that a valid reference sample contains at least one object, i.e. $\sum_{c\in C}N_{c}({\bm{x}})\geq 1$, where $C$ is the set of considered categories.
Counting hallucinations as deviations from the counting criteria. Given a generated sample $\hat{{\bm{x}}}$, we say $\hat{{\bm{x}}}$ exhibits a counting hallucination with respect to ${\mathcal{D}}_{\mathrm{\text{ref}}}$ if both of the following conditions hold:
  1. 1.
    Sufficient visual quality for counting. The image $\hat{{\bm{x}}}$ should exhibit sufficient visual quality such that the background and any present objects are clearly discernible. This ensures that counting hallucinations can be distinguished from other prominent failure modes, such as severe degradations. We define a binary indicator ${\mathbb{I}}_{\mathrm{CRI}}(\hat{{\bm{x}}})\in\{0,1\}$, referred to as the counting-ready indicator, where $1$ denotes that the image quality is sufficient for counting, and 0 indicates otherwise. In practice this indicator can be determined via human annotation, by a classifier calibrated against human annotations, or by other automated criteria.
  2. 2.
    Violation of counting facts. A violation of counting facts occurs if there exists a category $c$ such that $N_{c}(\hat{{\bm{x}}})\notin{\mathcal{S}}_{c}$, or if the image contains no objects (i.e., $\sum_{c\in C}N_{c}(\hat{{\bm{x}}})=0$) as the dataset requires at least one object per sample. In practice, we use a counting model to predict object counts in generated images as evidence for such violations.
A counting hallucination ($\mathrm{CH}$) is identified if its quality is sufficient for counting (${\mathbb{I}}_{\mathrm{CRI}}(\hat{{\bm{x}}})=1$) and the predicted counts deviate from the counting criteria ($\exists c:\,N_{c}(\hat{{\bm{x}}})\notin{\mathcal{S}}_{c}$ or $\sum_{c\in C}N_{c}(\hat{{\bm{x}}})=0$). Formally,
${\mathbb{I}}_{\mathrm{CH}}(\hat{{\bm{x}}})={\mathbb{I}}_{\mathrm{CRI}}(\hat{{\bm{x}}})\land\Big(\exists c:\,N_{c}(\hat{{\bm{x}}})\notin{\mathcal{S}}_{c}\;\lor\;\sum_{c\in C}N_{c}(\hat{{\bm{x}}})=0\Big).$
4.2 Datasets, Counting Hallucinations and Evaluation Protocol
The CountHalluset suite consists of three datasets for studying counting hallucinations: ToyShape, SimObject, and RealHand. Representative examples from each dataset are shown on the left side of Fig. 1. Additional statistical summaries are provided in Appendix A.4.
4.2.1 ToyShape
Dataset description. The ToyShape dataset comprises of 30,000 images with three geometric shapes: triangle, square, pentagon. All shapes are white, with an equal area of 120 pixels, and no shapes overlap within any image. The counting criteria specify that each image contains at most one instance of each shape category and at least one shape (in total), as shown in the top-left of Fig. 1. Formally, we have $C=\{\emph{triangle},\emph{square},\emph{pentagon}\}$. For each class $c\in C$, we set $N_{c}({\bm{x}})\in S_{c}=\{0,1\}$, subject to the constraint $\sum_{c\in C}N_{c}({\bm{x}})\geq 1$.
Counting hallucinations. According the counting criteria defined above, any image generated by the model trained on ToyShape that contains two or more shapes of the same category is considered a counting hallucination, such as an image containing two pentagons in the top-right region of Fig. 1. Besides, empty images (i.e., containing only the black background without any shapes) are also classified as counting hallucinations.
Counting model used in ToyShape. In our ToyShape generation experiments, we rely solely on a counting model to directly identify counting errors within images, without using a counting-ready indicator, as the dataset’s simplicity makes severely degraded images rare. We construct a large-scale ToyShape dataset of over 400,000 samples, with each category appearing 0–3 times per sample (i.e., $S_{c}=\{0,1,2,3\}$), to fine-tune a ResNet-50 (He et al., 2016). This setup captures a wide range of plausible scenarios, including normal samples and counting hallucinations, enabling accurate quantification. To further enhance the robustness of the counting model, we apply Gaussian noise during training. The model can achieve over 99.9% counting accuracy on generated data.
Quantification of counting hallucinations. Given a diffusion model trained on the ToyShape dataset, we generate a set of samples equal in size to the training dataset (30,000 images). We then employ the counting model to predict the number of objects in each generated image. Lastly, images that violate the counting criteria are identified as counting hallucinations, as illustrated in Fig. 2.
Figure 2: Evaluation procedure in the CountHalluSet suite. For ToyShape and SimObject, generated images are directly assessed for counting. For RealHand, a counting-ready indicator (a binary classifier aligned with human judgments) is introduced to separate counting hallucinations from other non-counting failures, such as severely deformed fingers. Count labels correspond to ToyShape: triangle, square, pentagon; SimObject: mug, apple, clock; RealHand: finger.
4.2.2 SimObject
Dataset description. The SimObject dataset includes 30,000 rendered images of everyday objects. It comprises three object categories: mug, apple, clock, each with 10 distinct intra-class variations. All objects are randomly placed on a wooden tabletop under fixed lighting conditions, with minimal mutual occlusion if any, as demonstrated in the middle of the left side of Fig. 1. The dataset is synthetically generated in Unreal Engine 5 (Epic Games, 2024), featuring photorealistic rendering with accurate simulation of object geometry, material attributes, lighting, and reflections. The counting criteria for this dataset are similar to those of the ToyShape dataset (see Sec. 4.2.1). Formally, we have $C=\{\emph{mug},\emph{apple},\emph{clock}\}$. For each class $c\in C$, we set $N_{c}({\bm{x}})\in S_{c}=\{0,1\}$, subject to the constraint $\sum_{c\in C}N_{c}({\bm{x}})\geq 1$.
Quantification of counting hallucinations. The definition of counting hallucinations, the construction of the counting model, and the quantification protocol follow the same procedure as in the ToyShape dataset, owing to the similarity of their counting criteria, as shown in Fig. 2.
4.2.3 RealHand
Dataset description. To investigate counting hallucinations in real-world scenarios, we construct the RealHand dataset comprising 5,050 human hand images sourced from the 11k Hands dataset (Afifi, 2019), Kaggle [^1] https://www.kaggle.com/, and Roboflow [^2] https://universe.roboflow.com/. Specifically, we aggregate raw images from these sources and remove those with blur or artifacts. Each image depicts a human hand with five fingers, capturing both dorsal (back) and palmar (front) views under varying lighting conditions and backgrounds (e.g., white surfaces, blackboards, walls), as shown in the lower-left of Fig. 1. The counting criteria specify that each image contains exactly five fingers, i.e., $C=\{\emph{finger}\}$ with $N_{c}({\bm{x}})\in S_{c}={5}$ for each $c\in C$, and $\sum_{c\in C}N_{c}({\bm{x}})=5$.
Counting hallucinations and non-counting failures. Generating realistic hand images is inherently prone to unexpected failures beyond counting hallucinations (e.g., extra or missing fingers), such as blurriness, severely deformed fingers, and visual artifacts. Therefore, we use a counting-ready indicator (i.e., a binary classifier) to identify these cases.
Counting-ready indicator and counting model used in RealHand. To separate counting hallucinations from previously mentioned failure modes, we finetune a MaxViT (Tu et al., 2022) model on 2.5k annotated generated images to serve as a counting-ready indicator aligned with human judgments. For counting, we finetune a YOLO-12 (Tian et al., 2025) model on 2k fingertip-annotated generated images. Based on human evaluation of 500 randomly sampled images by three annotators, the indicator and counting model achieve over 96% and 99% accuracy, respectively.
Quantification of counting hallucinations. Unlike the ToyShape and SimObject datasets, where all generated images are directly evaluated due to their simplicity and the rarity of severe failure cases (e.g., distorted or unrecognizable objects), the RealHand dataset requires an additional filtering step. Specifically, we first apply the counting-ready indicator to select hand images with sufficient visual quality, and only then use the counting model to predict the number of fingers. Figure 2 illustrates this distinction in detail.
5 Experiments
5.1 Implementation Details
This section outlines the experimental setup for both training and inference. All experiments are conducted on a cluster with 8 NVIDIA A100 GPUs for parallel training and inference.
Training of Diffusion Models. We train a DDPM (Ho et al., 2020) from scratch for the ToyShape dataset and fine-tune latent diffusion models (LDMs) (Rombach et al., 2022), pretrained on the CelebA-HQ dataset (Karras et al., 2018), for the SimObject and RealHand datasets. The time step $T$ is set to $1,000$ with a standard linear noise schedule. Images are resized to 128×128 for DDPMs and 256×256 and LDMs. We train the models using the Adam optimizer (Kingma, 2015) with a learning rate of 0.0001, running for 150k, 300k, and 80k steps on the ToyShape, SimObject, and RealHand datasets, respectively, with an effective batch size of 256.
Inference of Diffusion Models. For the ancestral sampling (i.e., DDPM), the sampling steps are set to 1,000, matching the diffusion steps during the training process. For the ODE-based solvers, we consider both first-order and second-order methods: DPM-Solver-1 (equivalent to DDIM (Song et al., 2021)) as a first-order solver, and DPM-Solver-2 (Lu et al., 2022) as a second-order solver. Each solver is evaluated under three widely-used denoising step configurations: 25, 50, and 100 steps. For the RealHand dataset, the counting model employs a YOLO detector with a confidence threshold of 0.3 and an IoU threshold of 0.1. Unless noted otherwise, the number of generated samples matches the number of training samples in each evaluation, and all experimental results are averaged over three different random seeds.
5.2 Quantifying Counting Hallucinations under Different Denoising Conditions
Table 1 presents quantitative results on counting hallucinations under varying denoising conditions, including solver types, ODE solver orders, and initial noise settings across three datasets: ToyShape, SimObject, and RealHand. Key observations are as follows.
(1) Increasing sampling steps in ODE solvers reduces counting hallucinations in synthetic datasets but exacerbates them in RealHand. As shown in Table 1, under commonly used ODE solver settings (25, 50, and 100 steps), increasing the number of sampling steps generally decreases the counting hallucination rate (CHR) in ToyShape and SimObject, but tends to increase it in RealHand, except at 100 steps with DPM-Solver-2. This contrasting behavior suggests that synthetic datasets benefit from finer solver granularity due to their simpler and more structured distributions, whereas real-world datasets such as RealHand exhibit more complex distributions, where additional steps may overfit local inconsistencies and thereby amplify hallucinations.
(2) Employing higher-order ODE solvers effectively mitigates non-counting failures but exacerbates counting hallucinations in RealHand. As shown in Table 1, comparing DPM-Solver-1 and DPM-Solver-2 on the RealHand dataset reveals that the latter reduces the non-counting failure rate but increases the counting hallucination rate. Nevertheless, this strategy consistently lowers the overall failure rates across datasets. For example, reducing CHR in synthetic datasets and TFR in RealHand, despite the increase in CHR observed in RealHand. These results underscore the need to develop ODE solvers that can jointly minimize non-counting failures while maintaining control over counting hallucinations in practical scenarios.
(3) DDPM substantially outperforms ODE solvers in mitigating both counting hallucinations and non-counting failures. Across all comparisons, ancestral sampling (i.e., DDPM) consistently achieves the lowest CHR, NCFR, and TFR, indicating that DDPM provides a practical lower bound for quantifying the failure rates of generative models. These results suggest that ancestral sampling serves as an effective strategy for minimizing hallucinations, particularly when computational efficiency is not the primary concern.
Table 1: Counting hallucination rate (CHR; the number of counting hallucinations divided by total generated samples) are reported under different solver configurations and initial noise settings across three datasets: ToyShape, SimObject, RealHand. For RealHand, we also report non-counting failure rate (NCFR) capturing non-counting failure samples. The total failure rate (TFR) is defined as the sum of counting hallucination rate (CHR) and non-counting failure rate (NCFR). Diffused and Normal refer to the ground-truth initial noise and standard Gaussian noise, respectively.
Solver Name
Sampling
Steps
CHR (ToyShape) CHR (SimObject) CHR (RealHand) NCFR (RealHand) TFR (RealHand)
Diffused Normal Diffused Normal Diffused Normal Diffused Normal Diffused Normal
DPM-Solver-1 (Song et al., 2021) 25 2.43 3.47 9.27 9.95 12.71 12.95 16.18 18.06 28.90 31.01
50 1.83 2.79 8.53 9.27 13.82 13.85 11.32 12.51 25.14 26.36
100 1.56 2.32 8.04 8.90 14.29 14.55 9.54 10.63 23.83 25.19
DPM-Solver-2 (Lu et al., 2022) 25 2.45 3.81 8.16 8.19 14.23 14.48 8.37 9.33 22.60 23.82
50 1.76 2.86 7.90 7.95 16.15 15.99 6.38 7.22 22.53 23.21
100 1.41 2.24 7.89 7.89 15.06 15.43 7.82 8.94 22.88 24.38
DDPM (Ho et al., 2020) 1000 0.63 0.64 4.98 4.99 10.82 10.75 2.07 2.39 12.88 13.14
(4) A better initial noise configuration can reduce both counting hallucinations and non-counting failures. As shown in Table 1, using “diffused” noise (i.e., ground-truth initial noise) tends to results in lower CHR, NCFR, and TFR compared to “normal” noise (i.e., standard Gaussian noise). This effect likely arises because diffused noise better aligns with the initial noise distribution encountered during training, suggesting that a more informed initialization can mitigate not only counting hallucinations but also other non-counting failures.
(5) Greater morphological complexity in countable objects increases the likelihood of counting hallucinations. Comparing the counting hallucination rates (CHR) across ToyShape, SimObject, and RealHand reveals that diffusion models are increasingly prone to counting hallucinations as object morphology becomes more complex, from simple geometric forms (e.g., triangle), to moderately complex natural objects (e.g., apple), to articulated biological structures (e.g., human finger). This trend arises because higher morphological complexity introduces intricate spatial relationships and fine-grained details that challenge the model’s ability to preserve accurate object counts. Consequently, the probability of erroneously adding or omitting object instances increases with structural complexity.
5.3 Correlation Studies: Counting Hallucination Rate vs. FID and Non-Counting Failure Rate vs. FID
This section examines the statistical correlations between the counting-based hallucination rate and the Fréchet Inception Distance (FID), as well as between the non-counting failure rate and FID, one of the most widely used evaluation metrics in image generation. Since counting hallucination represents a form of factualness hallucination, it is natural to ask: Does a better perception visual quality, indicated by a lower FID, necessarily imply fewer counting hallucinations by reflecting a smaller distribution gap between generated and training data? The answer is no necessarily. The following evidence supports this conclusion, with each paragraph presenting one piece of evidence.
The correlations between CHR and FID are dataset-dependent, rather than intrinsic. From Table 2, the Pearson correlation between counting hallucination rate (CHR) and FID is strongly positive in SimObject ($r$ > 0.87), indicating that lower FID corresponds to less counting hallucinations. In contrast, RealHand exhibits a strong negative correlation ($r$ = –0.9134), suggesting the opposite trend. Figure 3 directly illustrates this negative correlation.
The correlations between CHR and FID are solver-dependent, rather than intrinsic. As shown in Table 2, including DDPM (ancestral sampler) results reduces the consistency of the observed correlations between CHR and FID. For example, the Spearman coefficient $\rho$ on SimObject decreases from 0.8452 to 0.7538, whereas on RealHand it drops sharply from -0.8286 to -0.1429.
The correlations between NCFR and FID are consistently strong and positive across different solver types and sampling conditions. In contrast to the variable relationship observed between counting hallucination rate (CHR) and FID, the association between NCFR and FID is direct and robust. Specifically, whether or not DDPM results are included, both the Pearson ($r$) and Spearman ($\rho$) coefficients exceed 0.94, with very small p-values, indicating high statistical significance.
(a) Correlation trend on DPM-Solver-1.
(b) Correlation trend on DPM-Solver-2.
Figure 3: Correlation trends between FID and three failure rate metrics on RealHand, including counting hallucination rate (CHR), non-counting failure rate (NCFR), and totle failure rate (TFR), across different ODE solvers: (a) DPM-Solver-1 and (b) DPM-Solver-2.
Table 2: Correlation of counting hallucination rate (CHR), non-counting failure rate (NCFR), and total failure rate (TFR) with FID. ‘’incl.’ denotes DDPM results integrated with those from various ODE solvers across sampling conditions (e.g., initial noise, sampling steps, and solver orders).
CHR vs. FID NCFR vs. FID TFR vs. FID
Correlation measure SimObject RealHand RealHand RealHand
$r$ or $\rho$ p-value $r$ or $\rho$ p-value $r$ or $\rho$ p-value $r$ or $\rho$ p-value
Pearson 0.8762 0.0119 -0.9134 0.0109 0.9977 <<0.0001 0.9963 <0.0001
Pearson (incl. DDPM) 0.4925 0.1482 0.0602 0.8980 0.9902 0.0001 0.9017 0.0055
Spearman 0.8452 0.0041 -0.8286 0.0416 0.9429 0.0048 0.9999 <0.0001
Spearman (incl. DDPM) 0.7538 0.0118 -0.1429 0.7599 0.9643 0.0005 0.9999 <0.0001
5.4 How to Reduce Counting Hallucinations?
In recent years, significant improvements have been observed in the perceptual quality of images generated by diffusion models, as reflected in metrics such as FID. However, counting hallucinations remain prevalent. This may be because conventional metrics, including FID, primarily capture global perceptual realism and do not reflect semantic inconsistencies, such as incorrect object counts. As shown in Table 1 and the correlation analyses in Sec. 5.3, commonly applied numerical strategies, such as employing solvers with higher orders or increasing sampling steps, effectively reduce total failure rates but fail to address counting hallucinations in generating real-world images.
Interestingly, the lower counting hallucination rates observed in simpler countable objects (as shown in Observation 5 in 5.2) suggest that diffusion models can maintain accurate object counts when the underlying visual structure is sufficiently simple, such as in masked objects. This finding raises an important question: if diffusion models jointly denoise global visual features and factual constraints (e.g., object counts) within a shared latent space (which we refer to as joint-diffusion models (JDM)), could counting hallucinations be effectively reduced?
Specifically, we perform a channel-wise concatenation of the hand masks, which are obtained using SAM-2 (Ravi et al., 2024), to provide explicit structural constraints to the diffusion model. This mechanism directly steers the generative process at the pixel level, enforcing adherence to a predefined spatial layout. While the structural constraint is explicit, higher-level semantics, such as correct finger counts and plausible hand poses, emerge implicitly. By jointly learning visual representations and structural constraints within a shared latent space, the proposed JDM enforces semantic consistency and visual plausibility. As shown in Table 3, JDM consistently achieves lower failure rates compared to latent diffusion models (LDM), including CHR and NCFR, across diverse sampling conditions, highlighting its effectiveness in accurately generating structurally intricate objects that are susceptible to counting ambiguities.
Figure 4: Diagram of the proposed joint-diffusion model (JDM), which integrates hand segmentation masks as pixel-level structural constraints. By jointly learning visual representations and structural constraints within a shared latent space, JDM enforces high-level semantic consistency, leading to anatomically plausible hand poses and correct finger counts.
Table 3: Comparison of failure rates between the LDM and our proposed JDM on the RealHand dataset across different ODE solvers and sampling steps. Lower values indicate better performance.
Counting Hallucination Rate (CHR) $\downarrow$ Non-counting Failure Rate (NCFR) $\downarrow$
Solver Name DPM-Solver-1 DPM-Solver-2 DPM-Solver-1 DPM-Solver-2
Sampling Steps 25 50 100 25 50 100 25 50 100 25 50 100
LDM (Rombach et al., 2022) 12.95 13.85 14.55 14.48 15.99 15.43 18.06 12.51 10.63 9.33 7.22 8.94
JDM (ours) 11.41 10.67 10.51 10.12 9.66 10.07 2.94 2.42 2.29 3.51 4.63 2.26
6 Conclusion
In this work, we focus on a specific type of hallucination in diffusion models: counting hallucination. To systematically quantify this phenomenon, we introduce CountHalluSet, a dataset suite consisting of three datasets and a practical methodology for measuring counting hallucinations in image generation. Our quantification experiments yield several key findings. For example, commonly adopted numerical techniques in diffusion models help mitigate counting hallucinations in synthetic datasets but aggravate them in the RealHand dataset. Furthermore, our correlation analyses reveal that widely used perceptual quality metrics such as FID fail to capture the severity of counting hallucinations consistently.
Appendix A Appendix
A.1 Poof of the Bound of Total Accumulated Distributional Error in Ancestral Sampling
For the joint distributions of the ideal reverse process $q({\bm{x}}_{T:0})$ and the true reverse process $p_{\theta}({\bm{x}}_{T:0})$, their KL divergence is defined as:
${\mathcal{D}}_{\text{KL}}(q({\bm{x}}_{T:0})\|p_{\theta}({\bm{x}}_{T:0}))=\mathbb{E}_{q({\bm{x}}_{T:0})}\left[log\frac{q({\bm{x}}_{T:0})}{p_{\theta}({\bm{x}}_{T:0})}\right].$ (A.1)
For both reverse processes $q({\bm{x}}_{T:0})$ and $p_{\theta}({\bm{x}}_{T:0})$, we can decompose them as follows, leveraging the Markov property:
$q({\bm{x}}_{T:0})$ $=q({\bm{x}}_{T})\prod_{t=1}^{T}q({\bm{x}}_{t-1}\mid{\bm{x}}_{t}),$ (A.2)
$p_{\theta}({\bm{x}}_{T:0})$ $=p_{\theta}({\bm{x}}_{T})\prod_{t=1}^{T}p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t}).$ (A.3)
By applying the chain rule to decompose the KL divergence, we have:
${\mathcal{D}}_{\text{KL}}(q({\bm{x}}_{T:0})\|$ $p_{\theta}({\bm{x}}_{T:0}))=\mathbb{E}_{q({\bm{x}}_{T:0})}\left[log\frac{q({\bm{x}}_{T})\prod_{s=1}^{T}q({\bm{x}}_{s-1}\mid{\bm{x}}_{s})}{p_{\theta}({\bm{x}}_{T})\prod_{s=1}^{T}p_{\theta}({\bm{x}}_{s-1}\mid{\bm{x}}_{s})}\right]$
$=\mathbb{E}_{{q({\bm{x}}_{T:0})}}\left[log\frac{q({\bm{x}}_{T})}{p_{\theta}({\bm{x}}_{T})}+\sum_{t=1}^{T}log\frac{q({\bm{x}}_{t-1}\mid{\bm{x}}_{t})}{p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t})}\right]$
$=\mathbb{E}_{q({\bm{x}}_{T})}\left[log\frac{q({\bm{x}}_{T})}{p_{\theta}({\bm{x}}_{T})}\right]+\sum_{t=1}^{T}\mathbb{E}_{q({\bm{x}}_{t})}\left[\mathbb{E}_{q({\bm{x}}_{t-1}\mid{\bm{x}}_{t})}\left[log\frac{q({\bm{x}}_{t-1}\mid{\bm{x}}_{t})}{p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t})}\middle|{\bm{x}}_{t}\right]\right]$
$={\mathcal{D}}_{\text{KL}}(q({\bm{x}}_{T}\|p_{\theta}({\bm{x}}_{T}))+\sum_{t=1}^{T}\mathbb{E}_{q({\bm{x}}_{t})}[{\mathcal{D}}_{\text{KL}}(q({\bm{x}}_{t-1}\mid{\bm{x}}_{t})\|p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t}))].$ (A.4)
Then, according to the Data Processing Inequality (Cover & Thomas, 2012), marginalizing (or discarding some random variables) will not increase the KL divergence:
${\mathcal{D}}_{\text{KL}}(q({\bm{x}}_{0})\|p_{\theta}({\bm{x}}_{0}))\leq{\mathcal{D}}_{\text{KL}}(q({\bm{x}}_{T:0})\|p_{\theta}({\bm{x}}_{T:0})).$ (A.5)
The bound of total accumulated distributional error can be obtained by combining Eq. (A.1) and Eq. (A.5):
[…]
where equality (=) would strictly hold only in the ideal case where $q({\bm{x}}_{T})=p_{\theta}({\bm{x}}_{T})$ and $p_{\theta}({\bm{x}}_{t-1}\mid{\bm{x}}_{t})=q({\bm{x}}_{t-1}\mid{\bm{x}}_{t})$ for all $t$ and ${\bm{x}}_{t}$, which means the diffusion model can accurately recover the underlying data distribution $q({\bm{x}}_{0})$. This upper bound for the overall KL divergence (${\mathcal{D}}_{\text{KL}}(q({\bm{x}}_{0})\|p_{\theta}({\bm{x}}_{0}))$) between true data and model-generated data is identical to the variational bound derived in DDPM (Ho et al., 2020) when data entropy $H({\bm{x}}_{0})$ is excluded, thereby suggesting that optimizing DDPM’s variational loss directly tightens this explicit upper bound on the discrepancy between trued data and model-generated data.
Diffused-prior gap. The first source of errors arises from the mismatch between the distribution of the ground-truth diffused data $q_{T}({\bm{x}}_{T})=q_{0:T}({\bm{x}}_{T}|{\bm{x}}_{0})$ and the prior distribution used for sampling $p_{\theta}({\bm{x}}_{T})$. This mismatch stems from the fact that the number of diffusion steps $T$ is finite in practice. The discrepancy can be formally quantified using a general distributional metric:
${\mathcal{D}}_{\text{KL}}(q_{T}({\bm{x}}_{T})\|p_{\theta}({\bm{x}}_{T})),\quad p_{\theta}({\bm{x}}_{T})={\mathcal{N}}({\bm{0}},{\bm{I}}),$ (A.7)
where ${\mathcal{D}}_{\text{KL}}$ denotes the Kullback–Leibler (KL) divergence. Alternatively, other appropriate metrics such as FID or the Wasserstein distance can be employed to evaluate this distributional discrepancy. The diffused-prior gap, present from the outset, induces a deviation that persists throughout the denoising process, ultimately resulting in a sustained output error.
Noise model prediction error. The second source of errors stems from the noise prediction model. Because the model’s estimate ${\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)$ only approximates the true noise $\bm{\epsilon}$}, as guaranteed in principle by the Universal Approximation Theorem (Hornik et al., 1989; Nielsen, 2015; Yarotsky, 2022). Let $\Delta{\bm{\epsilon}}_{t}({\bm{x}}_{t})={\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)$ represent this prediction error. Based on Eq. (5), any discrepancy in ${\bm{\epsilon}}_{\theta}$ induces a corresponding deviation in the predicted mean at step $t$:
$\Delta{\bm{\mu}}_{t}({\bm{x}}_{t})=-\frac{\beta_{t}}{\sqrt{1-\beta_{t}}}\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\Delta{\bm{\epsilon}}_{t}({\bm{x}}_{t}).$ (A.8)
Since each reverse step uses the mean from the previous iteration, this bias accumulates throughout the denoising trajectory, ultimately contributing to the overall prediction error.
A.2 Poof of the Total Accumulated Trajectory Error in ODE-Based Sampling
Based on Eq. (7), we take a one-step numerical solver approximates the solution as example. Let ${\bm{x}}^{*}(\tau)$ denote the ideal trajectory that satisfies this ODE with the true score function $h^{*}({\bm{x}}(\tau),\tau)$ and starts from a ground-truth initial condition ${\bm{x}}^{*}(T)$ drawn from the true diffused distribution $q_{T}({\bm{x}}_{T})$. Similarly, $\hat{{\bm{x}}}_{k}$ represent the numerical trajectory computed by a numerical ODE solver at discrete time steps $\tau_{k}=T-k\Delta t$ for $k=0,1,\ldots,N$ and $\hat{h}$ is the estimated model function. For a specific numerical trajectory from $\hat{{\bm{x}}}_{0}-\hat{{\bm{x}}}_{N}$, we want to calculate its total accumulated trajectory error. Let $e_{k}=\hat{{\bm{x}}}_{k}-{\bm{x}}^{*}(\tau_{k})$ be the trajectory error at step $k$. Consider the transition from $\tau_{k}$ to $\tau_{k+1}=\tau_{k}-\Delta t$.
Ideal path evolution. Using the variation-of-constants formula for the exact solution over one step (from $\tau_{k}$ down to $\tau_{k+1}$):
${\bm{x}}^{*}(\tau_{k+1})={\mathcal{G}}(\tau_{k+1},\tau_{k}){\bm{x}}^{*}(\tau_{k})+\int_{\tau_{k}}^{\tau_{k+1}}{\mathcal{G}}(\tau_{k+1},u)h^{*}({\bm{x}}^{*}(u),u)\text{d}u,$ (A.9)
where ${\mathcal{G}}(t_{2},t_{1})=e^{\int_{t_{1}}^{t_{2}}f(v)\text{d}v}$ is the state transition operator for the linear part $\frac{\text{d}{\bm{z}}}{\text{d}\tau}=f(\tau){\bm{z}}$. Note that the integration limits are backward in time.
[…]
$\hat{{\bm{x}}}_{\text{exact}}(\tau_{k+1}\mid\hat{{\bm{x}}}_{k},\hat{h})={\mathcal{G}}(\tau_{k+1},\tau_{k})\hat{{\bm{x}}}_{k}+\int_{\tau_{k}}^{\tau_{k+1}}{\mathcal{G}}(\tau_{k+1},u)\hat{h}(\hat{{\bm{x}}}_{\text{exact}}(u\mid\hat{{\bm{x}}}_{k},\hat{h}),u)\text{d}u.$ (A.10)
The numerical method introduces a local truncation error $\delta_{k}$:
$\hat{{\bm{x}}}_{k+1}=\hat{{\bm{x}}}_{\text{exact}}(\tau_{k+1}\mid\hat{{\bm{x}}}_{k},\hat{h})+\delta_{k}.$ (A.11)
So, we have:
$\hat{{\bm{x}}}_{k+1}={\mathcal{G}}(\tau_{k+1},\tau_{k})\hat{{\bm{x}}}_{k}+\int_{\tau_{k}}^{\tau_{k+1}}{\mathcal{G}}(\tau_{k+1},u)\hat{h}(\hat{{\bm{x}}}_{\text{exact}}(u\mid\hat{{\bm{x}}}_{k},\hat{h}),u)\text{d}u+\delta_{k}.$ (A.12)
Error recurrence. Subtract the ideal evolution Eq. (A.9) from the numerical evolution Eq. (A.12):
$e_{k+1}=\hat{{\bm{x}}}_{k+1}-{\bm{x}}^{*}(\tau_{k+1})$
$={\mathcal{G}}(\tau_{k+1},\tau_{k})(\hat{{\bm{x}}}_{k}-{\bm{x}}^{*}(\tau_{k}))+\int_{\tau_{k}}^{\tau_{k+1}}{\mathcal{G}}(\tau_{k+1},u)\left[\hat{h}(\hat{{\bm{x}}}_{\text{exact}}(u\mid\hat{{\bm{x}}}_{k},\hat{h}))-h^{*}({\bm{x}}^{*}(u),u)\right]\text{d}u+\delta_{k}.$ (A.13)
Accumulated error formulation. We can unroll this recurrence relation from $k=0$ to N1N-1. Let ${\mathcal{G}}(\tau_{j},\tau_{k})={\mathcal{G}}(\tau_{j},\tau_{j-1}),\ldots,{\mathcal{G}}(\tau_{k+1},\tau_{k})$ for $j>k$, and ${\mathcal{G}}(\tau_{k},\tau_{k})={\bm{I}}$. Applying the recurrence repeatedly:
$e_{N}=$ ${\mathcal{G}}(\tau_{N},\tau_{0})e_{0}+\sum_{k=0}^{N-1}{\mathcal{G}}(\tau_{N},\tau_{k+1})$
$\cdot\left(\int_{\tau_{k}}^{\tau_{k+1}}{\mathcal{G}}(\tau_{k+1},u)\left[\hat{h}(\hat{{\bm{x}}}_{\text{exact}}(u\mid\hat{{\bm{x}}}_{k},\hat{h}))-h^{*}({\bm{x}}^{*}(u),u)\right]\text{d}u+\delta_{k}\right).$ (A.14)
Substituting $\tau_{N}=0$, $\tau_{0}=T$, $e_{N}=\hat{{\bm{x}}_{N}}-{\bm{x}}^{*}(0)$, and $e_{0}=\hat{{\bm{x}}}_{0}-{\bm{x}}^{*}(T)$:
$\hat{{\bm{x}}}_{N}-{\bm{x}}^{*}(0)=\underbrace{{\mathcal{G}}(0,T)(\hat{{\bm{x}}}_{0}-{\bm{x}}^{*}(T))}_{\text{propagated initial error}}+\sum_{k=0}^{N-1}{\mathcal{G}}(0,\tau_{k+1})$
[…]
where the propagated initial error term represents the deterministic impact on the final path error of a specific initial error instance, whose statistical origin lies in the distributional Diffused-Prior Gap.
A.3 The Discrepancy between DDPM and DDIM in the Forward and Reverse Process
As shown in Eq. (2), the forward process in DDPM (Ho et al., 2020) can be formulated as:
${\bm{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+(1-\bar{\alpha}_{t}){\bm{\epsilon}},\quad{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\bm{I}}).$ (A.16)
And the reverse process is defined as follows:
${\bm{x}}_{t-1}^{\text{DDPM}}=\frac{1}{\sqrt{\alpha_{t}}}{\bm{x}}_{t}-\frac{1-\alpha_{t}}{\sqrt{\alpha_{t}}\sqrt{1-\bar{\alpha}_{t}}}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)+\sqrt{\beta_{t}}{\bm{\epsilon}}.$ (A.17)
For DDIM (Song et al., 2021), the forward process is identical to that in Eq. (A.16), while its deterministic reverse process (with $\eta=0$) is defined as follows:
${\bm{x}}_{t-1}^{\text{DDIM}}$ $=\sqrt{\frac{\bar{\alpha}_{t-1}}{\bar{\alpha}_{t}}}({\bm{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t))+\sqrt{1-\bar{\alpha}_{t-1}}{\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t)$
$=\frac{1}{\sqrt{\alpha_{t}}}{\bm{x}}_{t}-\left(\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\alpha_{t}}}-\sqrt{1-\bar{\alpha}_{t-1}}\right){\bm{\epsilon}}_{\theta}({\bm{x}}_{t},t).$ (A.18)
Although DDPM and DDIM share the same forward (diffusion) process, DDIM does not simply strip the stochastic term from DDPM’s reverse (denoising) process. This is because the two methods employ different coefficient formulations for the predicted noise term.
A.4 Dataset Examples and Details
Figure A1 presents representative examples from the three datasets used in this study. As shown in Table A1, each dataset exhibits a balanced distribution across its respective categories. Table A2 further demonstrates that the distribution of images with different object counts is comparable, indicating the absence of object-count bias across datasets. The datasets appear in the same order in both tables, and the category order in Table A1 is defined as follows:
  • ToyShape: triangle, square, pentagon;
  • SimObject: mug, apple, clock;
  • RealHand: finger.
(a) Examples in the ToyShape dataset
(b) Examples in the SimObject dataset
(c) Examples in the RealHand dataset
Figure A1: Examples of datasets in the CountHalluSet suite: ToyShape, SimObject, and RealHand.
Table A1: Number of objects per category for each dataset.
ToyShape SimObject RealHand
Category 1 20,027 19,916 25,250
Category 2 19,902 20,101 -
Category 3 19,978 19,974 -
Table A2: Number of images by object count(s) per dataset.
ToyShape SimObject RealHand
1 object 10,003 10,022 -
2 objects 10,087 9,965 -
3 objects 9,910 10,013 -
5 objects - - 25,250
Total 30,000 30,000 25,250
A.5 Examples of Counting Hallucinations
We provide some representative examples of counting hallucinations corresponding to three datasets, as shown in Fig. A2. For the ToyShape and SimObject datasets, the counting hallucinations contain more than two instances of at least one category or do not contain any instances but with normal background, as shown in Fig. 2(a) and Fig. 2(b). In the Realhand dataset, counting-hallucinated samples are characterized by incorrect finger counts (e.g., 4 or 6 fingers), as shown in Fig. 2(c). Notably, the last two images of the first row in Fig. 2(c) appear to display five fingers; however, one finger is partially incomplete and visibly artifacted, effectively resulting in four valid fingers. Consequently, these samples are classified as counting hallucinations.
A.6 Examples of YOLO Predictions of Fingertips
Figure A3 presents representative examples of YOLO predictions on the RealHand dataset. The YOLO detector accurately identifies fingertips across diverse hand poses and orientations, including both dorsal and palmar views. However, it may still fail to recognize anatomically implausible configurations, such as hands with duplicated thumbs (e.g., the penultimate sub-figure of the first row and the last sub-figure of the last row in Fig. 3(a)) or hands with five middle fingers but without a thumb (e.g., the last sub-figure of the first row and the penultimate sub-figure of the last row in Fig. 3(a)). These abnormal samples are still detected as five-finger hands and therefore remain indistinguishable from anatomically valid ones in terms of finger counts.
Figure 3(b) shows several counting-hallucinated samples where YOLO predicts incorrect fingertip counts (e.g., 3, 4, 6, or 7 fingertips). Figure 3(c) further illustrates several rare edge cases (<1%) in which the detector fails to identify fingertips with unclear boundaries, cracked fingers, or floating artifacts. Such rare cases can be partly mitigated by tuning YOLO’s inference hyperparameters (e.g., IoU and confidence thresholds); however, doing so may compromise detection performance on the majority of normal samples. Since a consistent detection setup is applied across all evaluations on the RealHand dataset, the presence of these rare cases does not affect the validity of our overall conclusions, even though addressing them remains a challenging and low-reward effort.
A.7 Examples of Three Classes of Generated Hand Images
Figure A4 shows representative examples of three classes of generated hand images: counting-valid, counting-hallucinated, and non-counting failure samples. As illustrated in Fig. 2, the evaluation procedure of the RealHand dataset first applies the counting-ready indicator to separate counting-ready samples (i.e., counting-valid and counting-hallucinated) from images unsuitable for counting. The latter typically exhibit severe visual artifacts, such as distorted, cracked, partially visible, or floating fingers. Counting-ready samples are then further classified as counting-valid or counting-hallucinated based on whether the detected number of valid fingertips equals five.
(a) Examples of counting hallucinations corresponding to the ToyShape dataset
(b) Examples of counting hallucinations corresponding to the SimObject dataset
(c) Examples of counting hallucinations corresponding to the RealHand dataset
Figure A2: Examples of counting hallucinations corresponding to the ToyShape, SimObject, and RealHand datasets.
(a) Counting-valid YOLO predictions with five fingers
(b) Counting-hallucinated YOLO predictions with not five fingers
(c) Rare edge cases where YOLO predictions fail
Figure A3: Examples of YOLO predictions on generated RealHand images. (a) Counting-valid samples with five fingertip bounding boxes correctly predicted by YOLO (including cases with two thumbs). (b) Counting-hallucinated samples with incorrect fingertip count predictions (e.g., 3, 4, 6, or 7 detected fingertips). (c) Several rare edge cases where the counting-ready indicator fails to filter out and YOLO struggles to handle, such as blurred, incomplete, or anatomically disconnected fingers.
(a) Counting-valid samples (five valid fingers)
(b) Counting-hallucinated samples (not five valid fingers)
(c) Non-counting failure samples (unsuitable for counting)
Figure A4: Examples of three classes of generated hand images. (a) Counting-valid: images with five clearly discernible fingers. (b) Counting-hallucinated: images with an incorrect number of valid fingers (e.g., 4 or 6). (c) Non-counting failures: images unsuitable for counting due to severe visual artifacts, such as distorted, cracked, partially visible, or floating fingers.